Guardian Inc. Is trying to develop an asset-financing plan. The firm has $370,000 in temporary current assets and $270,000 In pemanent current assets. Guardian also has $470,000 in fixed assets. Assume a tax rate of 30 percent. a. Construct two alternative financing plans for Guardlan. One of the plans should be conservative, with 70 percent of assets financed by long-term sources, and the other should be aggresslve, with only 56.25 percent of assets financed by long-term sources. The current interest rate Is 12 percent on long-term funds and 8 percent on short-term financing. Compute the annual Interest payments under each plan. Annual Interest servelive Aggressive b. Given that Guardians earnings before interest and taxes are $250,000, calculate earnings after taxes for each of your alternatives. Earning After Taxes servalivs Aggressive c. What would the annual interest and eanings after taxes for the conservative and aggressive strategies be if the short-term and long-term interest rates were reversed? Conservative Aggressive Total interest Earnings after taxes
Answer #1

Please find below the answer. .please do provide rating..

Answer a)
Temporary current assets 370000
Permanent current assets    270000
Fixed assets                                               470000
Total assets 1110000
Conservative : Computation of annual interest
Interest rate Interest cost
Long term 12% =1110000*70%*12% 93240
Sort term 8% =1110000*30%*8% 26640
Total interest cost 119880
Aggressive : Computation of annual interest
Interest rate Interest cost
Long term 12% =1110000*56.25%*12% 74925
Sort term 8% =1110000*43.75%*8% 38850
Total interest cost 113775
Answer b) Computation of earning after tax
Conservative Aggressive
EBIT 250000 250000
Less Interest 119880 113775
EBT 130120 136225
Less Tax @ 30% 39036 40867.5
Earning after tax 91084 95357.5
Answer c)
Computation of interest cost if interest rate is reversed
Conservative : Computation of annual interest
Interest rate Interest cost
Long term 8% =1110000*70%*8% 62160
Sort term 12% =1110000*30%*12% 39960
Total interest cost 102120
Aggressive : Computation of annual interest
Interest rate Interest cost
Long term 8% =1110000*56.25%*8% 49950
Sort term 12% =1110000*43.75%*12% 58275
Total interest cost 108225
Computation of earning after tax
Conservative Aggressive
EBIT 250000 250000
Less Interest 102120 108225
EBT 147880 141775
Less Tax @ 30% 44364 42532.5
Earning after tax 103516 99242.5
Conservative Aggressive
Total interest 102,120.00                     108,225.00
Earning after tax 103,516.00                       99,242.50